# Joint-GCG: Unified Gradient-Based Poisoning Attacks on Retrieval-Augmented Generation Systems

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) systems enhance Large Language Models (LLMs) by retrieving relevant documents from external corpora before generating responses. This approach significantly expands LLM capabilities by leveraging vast, up-to-date external knowledge. However, this reliance on external knowledge makes RAG systems vulnerable to corpus poisoning attacks that manipulate generated outputs via poisoned document injection. Existing poisoning attack strategies typically treat the retrieval and generation stages as disjointed, limiting their effectiveness. We propose Joint-GCG, the first framework to unify gradient-based attacks across both retriever and generator models through three innovations: (1) Cross-Vocabulary Projection for aligning embedding spaces, (2) Gradient Tokenization Alignment for synchronizing token-level gradient signals, and (3) Adaptive Weighted Fusion for dynamically balancing attacking objectives. Evaluations demonstrate that Joint-GCG achieves at most 25% and an average of 5% higher attack success rate than previous methods across multiple retrievers and generators. While optimized under a white-box assumption, the generated poisons show unprecedented transferability to unseen models. Joint-GCG’s innovative unification of gradient-based attacks across retrieval and generation stages fundamentally reshapes our understanding of vulnerabilities within RAG systems. Our code is available at https://github.com/NicerWang/Joint-GCG.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) systems [1, 2] have emerged as a powerful paradigm for enhancing Large Language Models (LLMs). By coupling a retriever, which fetches relevant documents from an external corpus based on a given query, and a generator that synthesizes information to produce coherent and contextually appropriate responses, RAG systems leverage vast, up-to-date external knowledge. This architecture significantly improves the performance of diverse AI applications—including search engines [3], chatbots [4, 5], code assistants [6, 7], and knowledge bases [8, 9]—by ensuring outputs are accurate and up-to-date [10].
However, this remarkable power comes with a critical vulnerability: reliance on external corpora introduces the risk of corpus poisoning[11, 12, 13, 14, 15]. As illustrated in Figure 1, corpus poisoning involves malicious actors injecting crafted poisoned documents into the knowledge base. If retrieved and processed, these poisoned documents can cause the RAG system to generate wrong answers, harmful outputs, or biased opinions, thereby undermining its reliability. Besides, the growing trend of RAG systems employing open-source components is intended to facilitate transparency, customization, and data leakage prevention. However, this allows attackers to study and replicate system architectures meticulously, making RAG systems more susceptible to corpus poisoning attacks.
@@
Existing attack strategies, as detailed in Section 2, often adopt a fragmented approach, treating the retrieval and generation stages as disjoint optimization problems. For instance, Phantom [15] and LIAR [13] tackle the retriever and generator objectives independently and sequentially. Such methods can be suboptimal, as they overlook the synergistic effects that could be achieved by simultaneously optimizing for both components, potentially limiting the overall efficacy of the poisoning attack.
@@
2 Related Works
2.1 Retrieval-Augmented Generation (RAG) Systems
Retrieval-Augmented Generation (RAG) systems represent a significant advancement in mitigating the inherent limitations of Large Language Models (LLMs), such as knowledge cutoffs and hallucinations [1, 16]. The core principle of RAG is to ground LLM responses in external, verifiable knowledge. A typical RAG architecture involves a retriever and a generator. When a query is received, the retriever first fetches relevant documents or data snippets from a large external corpus (e.g., a vector database, enterprise knowledge base, or the internet) [17]. These retrieved contexts are then provided to the LLM (the generator) along with the original query. The LLM synthesizes this information to produce a more accurate, timely, and contextually appropriate response [16]. This reliance on external corpora, while beneficial for accuracy and currency, introduces new attack surfaces, particularly corpus poisoning, which is the focus of our work.
2.2 Adversarial Attacks on Large Language Models
Large Language Models, despite their impressive capabilities, are susceptible to various adversarial attacks [18]. These vulnerabilities include prompt injection, where malicious instructions are embedded in the input to hijack the model’s output [19], and training data poisoning of the base LLM itself, which can introduce subtle biases or backdoors [20]. These general vulnerabilities underscore the necessity for robust security measures throughout all stages of LLM development and deployment.
2.2.1 Gradient-Based Optimization for Adversarial Text Generation
Gradient-based optimization has become a cornerstone for crafting adversarial examples against LLMs. Early methods, such as HotFlip [21], utilized gradients related to input tokens to identify minimal character-level or token-level perturbations that could alter model predictions. Simple substitution techniques [22] and heuristic optimizations [23] have achieved only modest success, primarily with smaller models. In contrast, gradient-guided approaches have shown superior efficacy against robust transformer architectures. The Greedy Coordinate Gradient (GCG) attack [24] extended this by optimizing a universal adversarial suffix to elicit desired (often harmful) responses. Multi Coordinate Gradient (MCG), proposed in the Phantom work [15], further refines this by considering multiple substitutions simultaneously for efficiency. Improved versions, such as I-GCG [25], have also been explored. These methods demonstrate the power of gradient information but are primarily designed for direct attacks on the LLM, not the complex two-stage RAG pipeline.
2.3 Data Poisoning Attacks
2.3.1 Classical Data Poisoning in Machine Learning
Data poisoning is a type of malicious attack in which an attacker manipulates the training data of a machine learning model to compromise its behavior during inference [26]. By injecting a small quantity of carefully crafted malicious samples into the training dataset, attackers primarily seek to degrade overall model performance, cause misclassification for targeted inputs, or implant backdoors triggered by specific inputs [27, 28].
2.3.2 Attacks on Information Retrieval Systems
Information Retrieval (IR) systems have also been targets of manipulation. Traditional search engines faced "spamdexing" or "search engine poisoning," where malicious actors used techniques like keyword stuffing, hidden text, and link farms to artificially boost the ranking of certain web pages [29, 30]. Modern retrieval components, especially dense retrievers used in RAG, can be misled by imperceptible perturbations to documents, leading to ranking manipulation [31, 32]. Malicious attacks can force rankers to incorrectly order documents or retrieve irrelevant content, even with minimal modifications to the corpus.
2.3.3 Corpus Poisoning in RAG Systems
The reliance of RAG systems on external corpora makes them uniquely vulnerable to corpus poisoning. Existing research has explored various strategies: PoisonedRAG [11] focused on optimizing poisoned documents to maximize their retrieval probability. HijackRAG [33] applied similar principles to prompt leaking and spam generation. However, optimizing solely for retrieval can compromise the linguistic qualities needed to effectively steer the generator, often necessitating a larger number of injected documents. Sequential optimization approaches, such as applying HotFlip for retrieval and then GCG for generation, can be suboptimal because modifications made in one stage may negatively impact the other. LIAR [13] attempted to achieve better integration through an iterative loop but still lacks proper joint optimization, as it pre-assigns fixed optimizable lengths and optimization steps. Phantom [15] introduced trigger-based batch poisoning, but optimized retrieval and generation sequentially, which sometimes led to retrieval failures. TrojanRAG [14] targets the retriever model itself, a different threat model requiring greater access. In contrast, our Joint-GCG framework addresses these limitations by performing a truly joint optimization across both the retriever and the generator, marking a departure from these disjointed or sequential strategies.
@@ 3 Threat Model
Finally, this approach is consistent with methodologies in several contemporary RAG poisoning research works (e.g., LIAR[13], Phantom[15]), which also operate under white-box assumptions for both components, establishing this as a standard paradigm for in-depth analysis in this research area. The strong poison transferability demonstrated in our experiments (Section 5.2.5) reveals a practical path for relaxing this white-box assumption during attack deployment. An attacker can leverage Joint-GCG in a white-box setting using a powerful, locally hosted surrogate model that mimics the target’s possible architecture. The resulting poison can then be utilized to attack the target system, even if the components are different or entirely black-box. Our results show that poisons optimized on one generator can successfully attack others. Similarly, poisons demonstrate high transferability across different retriever architectures. This surrogate model approach transforms the attack into a gray-box scenario, where the attacker only needs to inject the pre-crafted document into the target corpus without internal access to the production models.
Gray-box Corpus Access: We assume the attacker has limited, or gray-box, access to the retrieval corpus. In our experiments, the attacker can inject a small, fixed number of poisoned documents into the corpus, typically one poisoned document per target query to ensure stealth, but cannot modify or delete existing legitimate documents. Our Adaptive Weighted Fusion (AWF) module calculates a stability metric by analyzing the top- documents retrieved for a query. Given that RAG systems often cite their sources, effectively making the top retrieved documents accessible, we find it realistic for the attacker to observe these results during the optimization phase. We also explore scenarios using synthetic corpora (Section 5.2.3) to mitigate this specific observation requirement. This gray-box corpus access reflects realistic scenarios such as decentralized knowledge bases, wikis, or systems that index publicly editable web content, where attackers can introduce malicious content but do not have control over the entire corpus. The constraint on the number of injected documents reflects the practical need for stealth, as injecting a large volume of suspicious documents would likely be detected.
@@ 4 Methodology: Joint-GCG Framework
Inspired by the success of Greedy Coordinate Gradient (GCG) techniques [24] in manipulating generator outputs through gradient-guided input optimization, Joint-GCG conceptualizes the RAG system as a single, integrated target for the poisoning attack, as illustrated in Figure 2.
@@
A core challenge in this unified approach lies in reconciling the different architectures, tokenization schemes, and vocabularies between the retriever and generator models. Our framework addresses these complexities through three key innovations: (1) Cross-Vocabulary Projection (CVP), which aligns the embedding spaces of disparate vocabularies; (2) Gradient Tokenization Alignment (GTA), which synchronizes token-level gradient signals across different tokenization outputs; and (3) Adaptive Weighted Fusion (AWF), which dynamically balances the attacking objectives for the retriever and generator. Together, these components enable a cohesive, gradient-based optimization strategy that jointly steers both the retrieval of the poisoned document and the subsequent generation of the desired malicious output.
4.1 Cross-Vocabulary Projection (CVP): Bridging Vocabulary Discrepancies
@@
The inherent vocabulary mismatch between retrievers and generators poses a significant challenge to joint gradient optimization. Typically, the retriever’s vocabulary differs from the generator’s vocabulary due to their distinct training corpora. CVP addresses this gap by utilizing a combination of the generator tokens to represent the retriever tokens in the embedding space.
@@
During the GCG-like attacks, an optimizable sequence was employed to steer the model output with gradient guidance. Let represent the optimizable sequence length and vocabulary size of the generator, and for the retriever. During the attack process, we compute gradient matrices and .
@@
Directly obtaining a high-dimensional linear transformation matrix using generator tokens to represent retriever tokens is rendered infeasible by both the ill-posed nature of constructing a sufficient system of equations and the substantial computational demands. Instead, CVP adopts a more tractable approach by focusing on mapping individual token embeddings. Let and represent the embedding matrices of the generator and the retriever, respectively, where denotes the embedding dimension. For each retriever token embedding , our objective is to find a representation such that it acts as a linear combination with the generator tokens when transformed by a learned embedding mapping function , closely approximates :
Here, represents a mapping function between the two embedding spaces.
CVP learns using shared tokens between the generator and retriever vocabularies, mapping their respective embeddings through an autoencoder. The autoencoder is trained to project the generator embedding of a shared token and then reconstruct the corresponding retriever embedding. This objective enables the autoencoder to learn a robust and semantically meaningful mapping that generalizes beyond the shared vocabulary, effectively capturing the underlying relationships between the two embedding spaces. The encoder of the trained autoencoder functions as , enabling us to solve a least-squares solution that transforms into and aligning the gradient information. Detailed architectural specifications of the autoencoder, training procedures, and analysis of CVP are provided in Appendix A.
4.2 Gradient Tokenization Alignment (GTA): Synchronizing Tokenization Granularity
@@
After CVP, another critical challenge arises from differing tokenization schemes. Retrievers and generators often employ distinct tokenizers, resulting in differences in how the input text is segmented into tokens. Consequently, the optimizable sequences for the retriever and generator may be tokenized into different sequences. This tokenization mismatch hinders direct gradient fusion. To address this, we propose GTA, a module designed to synchronize gradient signals at a finer, tokenizer-agnostic granularity.
@@
To achieve tokenization alignment, GTA employs character-level gradients as an intermediary. Retriever token gradients, derived from , are decomposed and assigned to their constituent characters, recognizing characters as a more fundamental and tokenization-agnostic text unit. Subsequently, retriever token gradients are constructed by averaging the gradients of the characters forming each generator token. This averaging method robustly consolidates character-level information back into the generator’s token space while mitigating potential noise from the decomposition process.
@@
Through the GTA process, we obtain a transformed retriever gradient matrix . Crucially, is now aligned with the generator’s gradient matrix in both vocabulary and sequence length dimensions. This alignment establishes a common gradient space, enabling a meaningful and direct fusion of gradient information from the retriever and the generator for joint optimization.
Additionally, we provide the pseudo-code of GTA in Appendix B.
4.3 Adaptive Weighted Fusion (AWF): Dynamically Balancing Retrieval and Generation Objectives
@@
With and now aligned in all dimensions, the final critical step is determining how to combine these gradient matrices to achieve joint optimization effectively. We propose AWF, a module that dynamically adjusts the relative contribution of each gradient matrix during the optimization process. This adaptive weighting mechanism is essential because the optimal balance between prioritizing retrieval and generation objectives can vary significantly depending on the specific attack scenario, the RAG system settings, and the characteristics of the target query and the corpus.
AWF fuses the gradients and weights the losses using a weighted sum controlled by an adaptive weighting parameter :
The weighting factor governs the relative influence of the retriever’s and generator’s gradient in the joint gradient update.
@@
As detailed in Appendix D, our experiments indicate a strong correlation between the retrieval rank of a poisoned document and attack success in RAG systems. We observed that poisoned documents retrieved at higher ranks are generally more influential in shaping the generator’s response. To optimize attack performance, it is desirable to position the poisoned document as high as possible in the retrieval ranking while ensuring a sufficient margin from other documents to mitigate potential rank fluctuations during subsequent optimization steps.
@@
AWF introduces a stability metric, , quantifying the robustness of the poisoned document’s retrieval rank:
@@
where and are the similarity scores of the query with the poisoned document and the benign document with the highest rank, respectively, is the average similarity score difference between consecutive documents, defined as below:
@@
This dynamic adjustment mechanism empowers Joint-GCG to adaptively balance the optimization of retrieval and generation objectives, resulting in more potent and robust poisoning attacks across diverse RAG system configurations.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) systems enhance Large Language Models (LLMs) by retrieving relevant documents from external corpora before generating responses. This approach significantly expands LLM capabilities by leveraging vast, up-to-date external knowledge. However, this reliance on external knowledge makes RAG systems vulnerable to corpus poisoning attacks that manipulate generated outputs via poisoned document injection. Existing poisoning attack strategies typically treat the retrieval and generation stages as disjointed, limiting their effectiveness. We propose Joint-GCG, the first framework to unify gradient-based attacks across both retriever and generator models through three innovations: (1) Cross-Vocabulary Projection for aligning embedding spaces, (2) Gradient Tokenization Alignment for synchronizing token-level gradient signals, and (3) Adaptive Weighted Fusion for dynamically balancing attacking objectives. Evaluations demonstrate that Joint-GCG achieves at most 25% and an average of 5% higher attack success rate than previous methods across multiple retrievers and generators. While optimized under a white-box assumption, the generated poisons show unprecedented transferability to unseen models. Joint-GCG’s innovative unification of gradient-based attacks across retrieval and generation stages fundamentally reshapes our understanding of vulnerabilities within RAG systems.
Code — https://github.com/NicerWang/Joint-GCG
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) systems (lewis2020retrieval; ram2023context) have emerged as a powerful paradigm for enhancing Large Language Models (LLMs). By coupling a retriever, which fetches relevant documents from an external corpus based on a given query, and a generator that synthesizes information to produce coherent and contextually appropriate responses, RAG systems leverage vast, up-to-date external knowledge. This architecture significantly improves the performance of diverse AI applications—including search engines (chatgptsearch), chatbots (vakayil2024rag; boulos2024nvidia), code assistants (zhou2022docprompting; nashid2023retrieval), and knowledge bases (siriwardhana2023improving; meduri2024efficient)—by ensuring outputs are accurate and up-to-date (fan2024survey).
However, this remarkable power comes with a critical vulnerability: reliance on external corpora introduces the risk of corpus poisoning (zou2024poisonedrag; xue2024badrag; tan2024glue; cheng2024trojanrag; chaudhari2024phantom). As illustrated in Figure 1, corpus poisoning involves malicious actors injecting crafted poisoned documents into the knowledge base. If retrieved and processed, these poisoned documents can cause the RAG system to generate wrong answers, harmful outputs, or biased opinions, thereby undermining its reliability. Besides, the growing trend of RAG systems employing open-source components is intended to facilitate transparency, customization, and data leakage prevention. However, this allows attackers to study and replicate system architectures meticulously, making RAG systems more susceptible to corpus poisoning attacks.
@@
Existing attack strategies, as detailed in the Appendix A, often adopt a fragmented approach, treating the retrieval and generation stages as disjoint optimization problems. For instance, Phantom (chaudhari2024phantom) and LIAR (tan2024glue) tackle the retriever and generator objectives independently and sequentially. Such methods can be suboptimal, as they overlook the synergistic effects that could be achieved by simultaneously optimizing for both components, potentially limiting the overall efficacy of the poisoning attack.
@@ 2 Threat Model
Finally, this approach is consistent with methodologies in several contemporary RAG poisoning research works (e.g., LIAR (tan2024glue), Phantom (chaudhari2024phantom)), which also operate under white-box assumptions for both components, establishing this as a standard paradigm for in-depth analysis in this research area. The strong poison transferability demonstrated in our experiments reveals a practical path for relaxing this white-box assumption during attack deployment. An attacker can leverage Joint-GCG in a white-box setting using a powerful, locally hosted surrogate model that mimics the target’s possible architecture. The resulting poison can then be utilized to attack the target system, even if the components are different or entirely black-box. Our results show that poisons optimized on one generator can successfully attack others. Similarly, poisons demonstrate high transferability across different retriever architectures. This surrogate model approach transforms the attack into a gray-box scenario, where the attacker only needs to inject the pre-crafted document into the target corpus without internal access to the production models.
Gray-box Corpus Access: We assume the attacker has limited, or gray-box, access to the retrieval corpus. In our experiments, the attacker can inject a small, fixed number of poisoned documents into the corpus, typically one poisoned document per target query to ensure stealth, but cannot modify or delete existing legitimate documents. Our Adaptive Weighted Fusion (AWF) module calculates a stability metric by analyzing the top- documents retrieved for a query. Given that RAG systems often cite their sources, effectively making the top retrieved documents accessible, we find it realistic for the attacker to observe these results during the optimization phase. We also explore scenarios using synthetic corpora to mitigate this specific observation requirement. This gray-box corpus access reflects realistic scenarios such as decentralized knowledge bases, wikis, or systems that index publicly editable web content, where attackers can introduce malicious content but do not have control over the entire corpus. The constraint on the number of injected documents reflects the practical need for stealth, as injecting a large volume of suspicious documents would likely be detected.
@@ 3 Methodology: Joint-GCG Framework
Inspired by the success of Greedy Coordinate Gradient (GCG) techniques (zou2023universal) in manipulating generator outputs through gradient-guided input optimization, Joint-GCG conceptualizes the RAG system as a single, integrated target for the poisoning attack, as illustrated in Figure 2. It adapts the iterative GCG process to optimize for a joint objective that harmonizes two distinct attack vectors: manipulating the generator’s output and ensuring the poisoned document is retrieved.
@@
To achieve this, we first define Joint Optimization Objective. This objective is minimized using an iterative optimization loop, which is similar to GCG. The primary challenge lies in combining gradients and losses from two architecturally disparate models. Joint-GCG addresses this through three key innovations: (1) Cross-Vocabulary Projection (CVP), which aligns the embedding spaces of disparate vocabularies; (2) Gradient Tokenization Alignment (GTA), which synchronizes token-level gradient signals across different tokenization outputs; and (3) Adaptive Weighted Fusion (AWF), which dynamically balances the attacking objectives for the retriever and generator.
3.1 The Joint Optimization Objective
As demonstrated in Figure 3, the attack is framed as an optimization problem to find an optimal adversarial sequence that, when injected into a document to form a poisoned document , minimizes a joint loss function for a given target query .
The weighting parameter is determined dynamically by AWF introduced in Section 3.2 to balance these two objectives.
Let be the set of Top-K documents retrieved from the corpus for the query , based on retriever model . Our loss function is defined as:
is an indicator function, which is 1 if is successfully retrieved into the Top-K context, and 0 otherwise. This formulation ensures that the generation loss is only considered when the poisoned document can actually influence the generator. The two loss functions are defined as:
Retrieval Loss (). To ensure is retrieved, we maximize its cosine similarity with the query by minimizing its negative value.
(3) -
Generation Loss (). To guide the generator towards a target answer , we use the cross-entropy loss to measure the difficulty of generating given the actual retrieved context .
3.2 The Joint-GCG Attack Loop
The entire Joint-GCG attack loop is designed to find a that minimizes . We now detail how each step of GCG algorithm is adapted for our joint attack task.
Step 1: Joint Gradient Computation and Fusion
The first step is to compute the gradient of with respect to the input tokens, which requires computing the gradients from the generator () and the retriever (). However, these raw gradients are incompatible due to differences in model vocabularies and tokenization schemes.
To resolve this, we employ Cross-Vocabulary Projection and Gradient Tokenization Alignment:
Cross-Vocabulary Projection (CVP): Bridging Vocabulary Discrepancies
@@ 3.2 The Joint-GCG Attack Loop
Typically, the retriever’s vocabulary differs from the generator’s vocabulary due to their distinct training process. CVP addresses this gap by utilizing a combination of the generator tokens to represent the retriever tokens in the embedding space.
@@
Let represent the token number of and the vocabulary size for the generator, and for the retriever. During the attack process, we compute gradient matrices and .
@@
Directly obtaining a high-dimensional linear transformation matrix using generator tokens to represent retriever tokens is rendered infeasible by both the ill-posed nature of constructing a sufficient system of equations and the substantial computational demands. Instead, CVP adopts a more tractable approach by focusing on mapping individual token embeddings. Let and represent the embedding layer matrices of the generator and the retriever, respectively, where denotes the embedding dimension. For each retriever token embedding , our objective is to find a representation such that it acts as a linear combination with the generator tokens when transformed by a learned embedding mapping function , closely approximates :
Here, represents a mapping function between the two embedding spaces. We use the encoder of a trained autoencoder to serve as the mapping function , by applying to the generator’s embeddings, we can then construct and solve a system of linear equations to obtain the final projection matrix enabling , via concatenating all solved . And then is used for transforming into . Detailed architectural specifications of the autoencoder, training procedures, and analysis of CVP are provided in Appendix B.1.
Gradient Tokenization Alignment (GTA)
@@
After CVP, another critical challenge arises from differing tokenization schemes. Retrievers and generators often employ distinct tokenizers, resulting in differences in how is segmented into tokens. To address this, we propose GTA, a module designed to synchronize gradient signals at a finer, tokenizer-agnostic granularity.
@@
To achieve tokenization alignment, GTA employs character-level gradients as an intermediary. Retriever token gradients are decomposed and assigned to their constituent characters, recognizing characters as a more fundamental and tokenization-agnostic text unit. Subsequently, to obtain token-level gradients for the retriever, we average the gradients of the characters within each token. This averaging method robustly consolidates character-level information back into the generator’s token space while mitigating potential noise from the decomposition process.
@@
Through the GTA process, we obtain a transformed retriever gradient matrix . Crucially, is now aligned with the generator’s gradient matrix in both sequence length and vocabulary dimensions, enabling a meaningful and direct fusion of gradient information from the retriever and the generator for joint optimization.
Adaptive Weighted Fusion (AWF): Fusing the Gradients
@@
With and now aligned in all dimensions, the final critical step is determining how to combine these gradient matrices to achieve joint optimization effectively. We propose AWF, a module that dynamically adjusts the relative contribution of each gradient matrix during the attack process. This adaptive weighting mechanism is essential because the optimal balance between prioritizing retrieval and generation objectives can vary significantly depending on the specific attack scenario, the RAG system settings, and the characteristics of the target query and the corpus.
@@
As shown in Appendix C.6, we observed that poisoned documents retrieved at higher ranks are generally more influential in shaping the generator’s response. To optimize attack performance, it is desirable to position the poisoned document as high as possible in the retrieval ranking while ensuring a sufficient margin from other documents to mitigate potential rank fluctuations during subsequent attack steps.
@@
AWF introduces a stability metric , quantifying the robustness of the poisoned document’s retrieval rank:
@@
where and are the similarity scores of the query with the poisoned document and the benign document with the highest rank, respectively, is the average similarity score difference between consecutive Top-K documents, defined as below:
@@
This dynamic adjustment mechanism empowers Joint-GCG to adaptively balance the optimization of retrieval and generation objectives, resulting in more potent and robust poisoning attacks across diverse situations.
Step 2: Candidate Set Generation And Re-evaluation
Using the joint gradient , we identify the most promising token substitutions. For each position in the adversarial sequence , we search for the Top- tokens from the vocabulary that would most likely decrease . Then we randomly select positions to perform the replacement and generate a set of candidate sequences . For each set in the set, we perform a forward pass involving recalculating both and , then calculate for all candidates using Equation 2.
Step 3: Greedy Selection and Update
Finally, we perform a greedy selection. From all evaluated candidates, we choose the one that yields the lowest and update accordingly for the next iteration.
After multiple optimization iterations, we can achieve the Joint Optimization Objective in Eq 1, and simultaneously ensure the poisoned document is retrieved and the generator’s output is manipulated.